OpenStudy (anonymous):

C. Write an equation of the perpendicular to y = 3x + 1 passing through (0, -4)

OpenStudy (stefrheart):

First you need to do the opposite reciprocal of this slope to get the perpendicular slope

OpenStudy (anonymous):

i have no clue what any one this means im in a slow math and im behind can you help me start it?

OpenStudy (stefrheart):

No problem! :) its okay lets start with the slope of this formula

OpenStudy (anonymous):

okay

OpenStudy (stefrheart):

What is the slope?

OpenStudy (anonymous):

3?

OpenStudy (stefrheart):

Yes! Now the what is the opposite of 3?

OpenStudy (anonymous):

1?

OpenStudy (stefrheart):

Not really. The opposite would be literally the opposite. For example: The opposite of 1 is -1 so what is the opposite of 3?

OpenStudy (anonymous):

oh -3

OpenStudy (stefrheart):

Great! Now the reciprocal is a tad more difficult For example: The reciprocal of 10 is 1/10 so what is the reciprocal of -3?

OpenStudy (anonymous):

1/-3 0r 1/3?

OpenStudy (stefrheart):

It would be -1/3 because we did the reciprocal of -3

OpenStudy (anonymous):

okay,

OpenStudy (stefrheart):

Do you know point slope formula?

OpenStudy (anonymous):

no

OpenStudy (stefrheart):

Its y-y1=m(x-x1)

OpenStudy (stefrheart):

Do you know what to do next?

OpenStudy (anonymous):

okay, so not really,

OpenStudy (stefrheart):

Well the y1 represents the y coordinate in the given coordinate and the x1 represents the x coordinate in the given coordinate what is your given coordinate?

OpenStudy (anonymous):

(0,-4)

OpenStudy (stefrheart):

Now which is x1 and which is y1?

OpenStudy (anonymous):

-4 x? y 0?

OpenStudy (stefrheart):

Yes! So now substitute the x1, y1, and m

OpenStudy (stefrheart):

On y-y1=m(x-x1)

OpenStudy (anonymous):

do you have a facebook because openstudy keeps freezing?

OpenStudy (stefrheart):

no i dont sorry

OpenStudy (stefrheart):

Here this is what it would look like y-0=-1/3(x+4)

OpenStudy (anonymous):

that's the full equation?

OpenStudy (stefrheart):

no now you need to simplify it

OpenStudy (anonymous):

okay, what part would I start with?

OpenStudy (stefrheart):

Well the 0

OpenStudy (anonymous):

how do I start it tho like whats the first step to simplify

OpenStudy (stefrheart):

Well first the 0 goes away since its 0 so it becomes y=-1/3(x+4)

OpenStudy (anonymous):

okay so next would we do x+4? and it would be 4x?

OpenStudy (stefrheart):

No you would have to distribute the -1/3

OpenStudy (anonymous):

I don't know how to do that..

OpenStudy (stefrheart):

You never learned how to distribute?

OpenStudy (anonymous):

I do cyber school and the teachers blow throw things and don't review them that's why im behind

OpenStudy (stefrheart):

Oh its okay this is how it would look like: 4(x+4) now you multiply the x by 4 and the 4 times 4 so it looks like 4x+16

OpenStudy (anonymous):

okay so the equation would look like y=-1/3(4x+16)?

OpenStudy (stefrheart):

no i was using the 4x+16 as an example

OpenStudy (stefrheart):

you need to multiply the -1/3 to x and -1/3 to 4

OpenStudy (anonymous):

okay so -1/*3 times 4 is -1.333

OpenStudy (stefrheart):

or you could leave it in fraction form

OpenStudy (anonymous):

im so confused the way im doing it the decimal is -1.3333333333 and so on, and then the fraction form is -4/3

OpenStudy (stefrheart):

yes most teachers like it in fraction form so now what is the equation?

OpenStudy (anonymous):

y=-1/3(x+-4/3)

OpenStudy (stefrheart):

well close! its y=-1/3x-4/3